A charge Q is brocken in two parts and placed at some distance.Find that the force between them is maximum.

Grade:12

1 Answers

Arun
6 years ago
 

let parts are sepereted at a distance of d

 

so repulsion force F= k q(Q-q)/d2

 

     here Q and d are constant so for maximum force differentiate the force w.r.t q

 

 dF/dq = k/d[Q -2q]

 

 for maximum force dF/dq=0

 

                             Q-2q=0

 

                              Q/q =2